What is the phase shift of y= 3- 4 cos pi/2 (2x+ 1/2)
Answer by lwsshak3(11628)

What is the phase shift of y= 3- 4 cos pi/2 (2x+ 1/2)
***
form of equation for cos function:
y=Acos(Bx-C)+d, A=amplitude,Period=2π/B,Phase shift=C/B,d=vertical shift
For given equation:
Rewrite:
y=-4cos(π/2*2x+π/2*1/2)+3
y=-4cos(πx+π/4)+3
B=π
C=π/4
phase shift=C/B=(π/4)/π=1/4
phase shift=1/4 (shift to the left)